• Rome Lab Weather Application

– Goal is to graphically represent the weather using VRML ontop of the VRML terrain data.

– We are trying to predict Lake effect snow.

• People:

– Scott Klasky (20 hours / week) (Project Leader)

– Hon Yau (0? hours / week)

• Recent Accomplishments (NOT MUCH)

– Jan: Decided to use the Advanced Regional Prediction System from the Center for Analysis and Prediction of Storms (CAPS)

– Feb 28: NCAR graphics was installed on our systems.

– Feb 29: Code was run using examples.

• Key Milestones/Deliverables/Plan – Prediction of Lake effect snow from code (May 1, 1996).

– VRML output (July 1, 1996).

– Incorporation of VRML output in terrain rendering (August 1)

– Fancier output in VRML (September 1, 1996)

• Constraints/Problems/Inhibitors

1. Need more contact with CAPS people. I am in the process of trying to set up a meeting. Until this can take place, not much can be done.

2. Need someone else on this project.

3. Need initial data for Lake Effect Snow. We will use Feb, 14, 1996. Bob Farallel, the Meteorologist at Rome Labs, is trying to get us this data set.